katherine:

Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form. y = x + 8 y = x − 8 y = −5x − 10 y = −5x + 20

katherine:

@563blackghost

563blackghost:

Find the slope first. \(\Large\bf{\frac{10-5}{2+3}}\)

katherine:

5/5

katherine:

next?

katherine:

@Ultrilliam

563blackghost:

sorry i was afk.... Correct! So that means it has a slope of 1. \(\large\bf{y=x+b}\) We plug a point to find b. \(\large\bf{5=-3-b}\) what is your equation?

katherine:

A?

katherine:

idek,sorry x'D

563blackghost:

Yea A :D

katherine:

80!! TYSMMM

563blackghost:

your welcome :D